OpenStudy (anonymous):

how do you factor 81a^2-100b^6

hartnn (hartnn):

81 is the square of ?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

\[81a^2-100b^2\]

hartnn (hartnn):

do u know the square root of 81 and 100 ?

OpenStudy (anonymous):

9 & 10

hartnn (hartnn):

right. so u can write that as \(\large (9a)^2-(10b)^2\) ok?

hartnn (hartnn):

now use \(\huge a^2-b^2=(a+b)(a-b)\)

OpenStudy (anonymous):

i understood the first part you lost me at the second message

OpenStudy (anonymous):

& being that its \[100b^6\] would it be \[( 10b)^6\]

hartnn (hartnn):

is it 100b^6 because you wrote it as 100b^2 afterwards

OpenStudy (anonymous):

oh okay , right

hartnn (hartnn):

so what is it?

OpenStudy (anonymous):

\[100b^6\]

hartnn (hartnn):

ok, so 100 b^6 can be written as (10b^3)^2 ok?

OpenStudy (anonymous):

oh okay because you multiply the exponents?

hartnn (hartnn):

yup.

hartnn (hartnn):

now u have (9a)^2 - (10b^3)^2 use the basic identity A^2-B^2 to factor this: put A=9a, B = 10b^3

hartnn (hartnn):

\(\huge A^2-B^2=(A+B)(A-B)\)

hartnn (hartnn):

so (9a)^2- (10b^3)^2 will be (9a+10b^3)(9a-10b^3) did u understand ?

OpenStudy (anonymous):

not at all :(

hartnn (hartnn):

which part ? upto here (9a)^2- (10b^3)^2?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

now there is a basic identity to be used : \(\huge A^2-B^2=(A+B)(A-B)\) put A=9a and B=10b^3 here

OpenStudy (anonymous):

ok

hartnn (hartnn):

did u get (9a+10b^3)(9a-10b^3) ?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

thats your final answer then :)

OpenStudy (anonymous):

^_^ Really?

hartnn (hartnn):

YES!! absolutely :)

hartnn (hartnn):

u have factored that successfully

hartnn (hartnn):

factoring means bringing in the form of (...)(...)

OpenStudy (anonymous):

okay i just tried the next one can you correct me if im wrong

hartnn (hartnn):

sure.

OpenStudy (anonymous):

\[64a^4-25y^4\] thats the question .

OpenStudy (anonymous):

\[(8a^2)^2-(5y^2)^2\]

hartnn (hartnn):

that is correct, and what will u do to factorize ?

OpenStudy (anonymous):

that idk

hartnn (hartnn):

use same thing A^2-B^2 = ??

OpenStudy (anonymous):

(a+b)(a-b)?

hartnn (hartnn):

yup.so what will be \((8a^2)^2-(5y^2)^2\)

OpenStudy (anonymous):

that system works for them all ?

hartnn (hartnn):

yes, for anything of the form of difference of squares

OpenStudy (anonymous):

aslo , if one of the numbers arent a perfect square its unfactorable ?

hartnn (hartnn):

then check whether its a perfect cube....but it really depends on problem